OpenStudy (anonymous):

I will medal and fan! An equation is shown below: 9(3x - 16) + 15 = 6x - 24 Part A: Write the steps you will use to solve the equation, and explain each step. Part B: What value of x makes the equation true?

rvc (rvc):

open the brackets

OpenStudy (anonymous):

so it is 27x and 144

OpenStudy (anonymous):

right?

OpenStudy (anonymous):

are you there?

rvc (rvc):

u have \[\Large 9(3x - 16) + 15 = 6x - 24 \] \[\Large 9(3x)-9(16)+15=6x-24\]

rvc (rvc):

whats 9(3x)=? 9(16)=?

OpenStudy (anonymous):

27x and 144

rvc (rvc):

correct \[\Large 27x-144+15=6x-24\]

rvc (rvc):

Now bring the like terms together

OpenStudy (anonymous):

that part is confusing i don't know how to do it

rvc (rvc):

keep all the terms with the x on the LHS and the constants on RHS

rvc (rvc):

\[\huge 5 apples - 5 = 12 apples +6\] \[\huge 5apples-12apples=6+5\]

rvc (rvc):

u here?

OpenStudy (anonymous):

yes then what?

rvc (rvc):

did u keep the like terms together?

rvc (rvc):

may i see the work?

OpenStudy (anonymous):

i don't get it can you just show me so i can understand it?

rvc (rvc):

\[\huge\color{red}{27x-6x}=\color{blue}{144-15-24}\]

rvc (rvc):

now solve further

OpenStudy (anonymous):

what is that supposed to mean?

rvc (rvc):

144-15=?

OpenStudy (anonymous):

129

rvc (rvc):

129-24=?

OpenStudy (anonymous):

105

rvc (rvc):

27x-6x=?

OpenStudy (anonymous):

21x

rvc (rvc):

yeah \[\Huge \rm Finally~equation~reduces~to\] \[\Huge\color{gold}{ \bf 21x=105}\]

rvc (rvc):

x=?

OpenStudy (anonymous):

5

rvc (rvc):

correct! All the best!

rvc (rvc):

\[\huge\color{purple}{\frak \bigstar Welcome~to~openstudy~\bigstar}\]